Graph using elimination
2x+6y=8
3x+9y=12

Answer by stanbon(75887)

Solve using elimination
2x+6y = 8
3x+9y =12
-------
Modify::
x + 3y = 4
x + 3y = 7
-----
Subtract:
0 = 3
----
That is a contradiction which resulted from
assuming there were one or more points of the form (x,y).
------
That is a wrong assumption; there is no such solution.
------------------
Note: the graphs of the two equations are parallel.
Cheers,
Stan H.
============